OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 1) and (3, 4)

OpenStudy (saifoo.khan):

m = 3/2

OpenStudy (saifoo.khan):

c = -3/2

OpenStudy (saifoo.khan):

y = 3x/2 - 3/2 2y = 3x - 3